2017/3/13(Tue.)

 * Remote participants by eZuce: None

==== Report of work (*:responsible person) ====

 * Alignment, ETMY: *Enomoto, Miyo, Miyamoto
  * Diagonal of Bottom filter matrix.
  * Feedforward using accelerometer to IP error point.
  * Alignnemtn for IFO will be checked every night.
  * IMC was aligned last night, but MCi had ~100urad shift. MC alignment will be check every night.

 * Camera: Yokozawa
  * Work of IX, IY inside chambers were done. Pump down will be start tomorrow.
  * Team will start this week.

 * IMMT: Koseyama
  * IMMT had jumps several times but it is unknown whether other SUS had.
  * It is not clear that the duct connection or kicking table.
  * Oplev will be centered today.

 * CRY: Mirror reached at 16K.


==== Work plan (*:responsible person) ====

 * VAC: Uchiyama
  * Duct connection, but they will wait for centering of oplev of IMMT.
 * Tcam: 
  * phi800 door of IX and IY will be closed today.
  * Pump down will be start tomorrow. 

 * Alignment, ETMY: *Enomoto, Miyo, Miyamoto
  * MC alignment will be check every night.
  * Alignnemtn for IFO will be checked every night.


 * IMMT: Koseyama
  * Oplev will be centered today.



==== Laser hazard ====
 * [AM] IR: PSL, Green: off.
 * [PM] IR: PSL, Green: off. 


==== Weekly ====
 * week of 3/12
  * 3/14: meeting for pump-down check
  * TMS installation at ETMY.
  * Connecting ducts at the corner area.

 * week of 3/19
  * Vacuum test at at the corner area.
  * TMS alignment using green light at ETMY.
  * Beam alignment to REFL port.


 * Issues
  * Request of using light in the night from Alignment and calibration.
  * Connection of PD signal to DAQ should be implemented this week or next week.
  * GigE camera should be placed at the REFL port eventually.